OpenStudy (akonkel):

Micah ate 5 more pieces of pizza than Emily. Together they ate 17 pieces of pizza. How many pieces did Micah eat?

OpenStudy (akonkel):

@Godlovesme @jagr2713

jagr2713 (jagr2713):

no more please

OpenStudy (akonkel):

well actually one more after this

OpenStudy (danjs):

The information gives you two equations Total pieces = Micah + Emily and Micah + 5 = Emily

jagr2713 (jagr2713):

yea i cant b.c its to easy lol

OpenStudy (danjs):

17 = M + E and E = M + 5 solve for E and M

OpenStudy (akonkel):

lol wow well i suck at math what grade r u

jagr2713 (jagr2713):

12

jagr2713 (jagr2713):

Dan knows everything so listen to him

OpenStudy (akonkel):

i dont have 12 as an option

OpenStudy (akonkel):

i have 17 though

jagr2713 (jagr2713):

huh

OpenStudy (akonkel):

tthese r my options 17 5 11 6 8.5

OpenStudy (danjs):

17 = M + E and E = M + 5 youw ant to know M Sub in E = M + 5 into the first equation 17 = M + M + 5 17 = 2M + 5 12 = 2M M = 6 pieces

OpenStudy (akonkel):

so what would my answer be ?

OpenStudy (akonkel):

6 or 17

OpenStudy (danjs):

Micah ate 6 pieces Emily ate 11 pieces Total together = 17 pieces

OpenStudy (akonkel):

@jagr2713 give danjs a medal he earns it and u do to for trying hard thanx both!!! :)

jagr2713 (jagr2713):

thx

OpenStudy (akonkel):

yup :)